Wednesday, September 18, 2013

Lost Balls - If Found, Please Return to the Blue CCD...

Late last night we considered the grim possibility that our problems may be due to the blue CCD being tilted...  A quick look back at the photos taken before we attached the dewars confirmed this suspicion: Aaaaarrrrgggh!!!  

Before going in to inspect the damage this morning, we got to celebrate Vic's birthday :)  HB2U - HB2U - HB2Vvvv - HB2Uuuuuu! :)


 Thanks for the fortifying snacks - the sugar buzz helped us deal with what was to follow...

David bravely suited up & attached his grounding strap.

Then set about opening up the various hatches situated around the blue dewar to see if what we suspected was indeed the case. 

It's a long story, but in short: the red camera had to be taken apart in Durham to fix various problems.  Along the way, the guys found that the sawn-off ball bearings attached to the 3 micrometers that provide tip/tilt adjustability for the plate supporting the CCD were loose.  They decided to bond the balls in place & all was well.  The blue cryostat did not need to be opened up to fix anything so the risk associated with exposing such delicate bits for the sake of securing the balls was not justified.  But now, seeing the chip misaligned, the fear was that the blue balls may have worked themselves out of position.  So the micrometer access hatches had to be opened to take a look inside...

Micrometer #1 looked ok - you can see the ball on the tip of the micrometer seated nicely in the pocket in the plate on the left.

Not so for the other 2 though!  The balls had indeed come off the ends of the micrometers & fallen down into the dewar :(  Grrrrr!!!!

Thus wrecking the tip/tilt of the plate that supports the CCD...


So then the trick was to hunt down & extract the free-ranging ball bearings.  Sherlock Bramall on the case!

Ah Hah!   There's one of the little buggers - lurking in a dark corner...

 Soon to be extracted with the aid of a kapton-coated allen key & a pair of tweezers.

Mugshots of 1 of the criminals - seen from below & above.

So the next challenge would be to put them back into place, somehow.  A crafty plan was devised which involved bonding the balls to long bolts that could reach into the inaccessible spaces.  Once securely positioned, the bolts will be snapped off, hopefully leaving the balls where they belong.  The epoxy will take a while to cure sufficiently so we'll tackle this tomorrow morning.

Meanwhile, up on the tracker Janus & Eben removed the FIF which needs to come back down for some lab testing.  We want to peak-up the FIF stage positions using the actual HRS fibres so that we can offset to each fibre position & know that we're optimally directing the starlight into the fibre.

We also need to do some throughput tests before we route the fibre cable up to the payload so we're adapting Jürgen's fibre test rig for the task.

Back in the clean room we suited up journalist Linda & Luke (yes, we have another 1!) for a tour of the HRS.

Chief tour operator Dr Schmoll in action again!  We look forward to reading all about SALT HRS in Nature News in a few weeks time :)

David Buckley came up today & brought along Luke's beloved & much-missed bike.  No more excuses now - time to get back into training during all that spare time between midnight & 8am!

Sanity's finally returned to the control room with just Fred & Brent working their way through a clear, stable night with good seeing.  They almost look a bit lonely like this...

No comments:

Post a Comment